OpenStudy (anonymous):

What is the slope of the line 2x - 5y = -10?

OpenStudy (anonymous):

The standard form of the slope - intercept form is y = mx + c, where 'm' is slope. So first we need to convert given equation into standard form 2x - 5y = -10 -5y = -10 - 2x y = 2 + (2/5)x. y =(2/5)x + 2 So slope is m = 2/5..